achieve a hwav n volts of potential ifference must an electron, initilly a rst, be accelerated to achieve a wavelength of 0.27nm?

please write neatly
Answer #1

From eV = hc/lemda

V = hc/e. Lemda

V = 1240 eV.nm /e. 0.27nm

V = 1240/0.27 = 4592.6 Volt